OpenStudy (anonymous):

Represent \[\sin^{-1} \] (2ax sqrt{1-a^2 x^2}), -1/sqrt{2} le ax ge -1/sqrt{2} in the simplest form.

OpenStudy (anonymous):

\[\sin^{-1} (2ax \sqrt{1-a^2x^2}), \]\[-1/\sqrt{2} \le 1/\sqrt{2} \] in the simplest form

myininaya (myininaya):

\[ \sin^{-1}(2ax \sqrt{1-a^2 x^2}), -1/\sqrt{2} \le ax \ge -1/\sqrt{2} \]

myininaya (myininaya):

lol

OpenStudy (dumbcow):

hmm this looks ugly :)

OpenStudy (anonymous):

sorry i meant greater than or equal to 1/sqrt{2}

OpenStudy (dumbcow):

im not sure what to do really

myininaya (myininaya):

i kindof don't understand the question

OpenStudy (anonymous):

what to do now

myininaya (myininaya):

so this is exactly how the questions reads?

myininaya (myininaya):

?

OpenStudy (dumbcow):

is it \[-1/\sqrt{2}\le ax \le 1/\sqrt{2}\]

myininaya (myininaya):

yea the inequality you had didn't make sense

myininaya (myininaya):

say something back please lol

OpenStudy (dumbcow):

so \[\pi/2 \le \sin^{-1}(2ax \sqrt{1-a^{2}x^{2}}) \le 3\pi/2\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

what?

myininaya (myininaya):

how did you get that cow?

OpenStudy (anonymous):

yes how did u get that dumbcow

OpenStudy (dumbcow):

i plug in 1/sqrt2 and -1/sqrt2 for ax

myininaya (myininaya):

so thats what we want?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok...

OpenStudy (dumbcow):

really thats it? bad question wording that was not simplifying

OpenStudy (anonymous):

hmn

OpenStudy (anonymous):

so what can i do now 1

OpenStudy (anonymous):

no help

myininaya (myininaya):

i don't know what the question is sorry

OpenStudy (anonymous):

r u on facebook

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

PatrickJMT: making FREE and hopefully useful math videos for the world!